\begin{thebibliography}{999}
\bibitem{BSR99}
{Bigatti, {A.M.}; La Scala, R.; Robbiano, L.} {Computing toric ideals}. {\em J. Symb. Comput.} \textbf{27} (1999) 351--365.

\bibitem{BR15}
{Boocher, A.; Robeva, E.} {Robust toric ideals}. {\em J. Symb. Comput.} \textbf{68}(1) (2015), 254--264.

\bibitem{BCO1}
{Branco, M.B.; Colaço, I., Ojeda, I.} {Generalized repunit numerical semigroups} (work in progress).

\bibitem{singular}
{Decker, W.; Greuel, G.-M.; Pfister, G.; Sch{\"o}nemann, H.}
{{\sc Singular} {4-2-1} --- {A} computer algebra system for polynomial computations.} \url{http://www.singular.uni-kl.de} (accessed on 28 October 2021).

\bibitem{Philippe}
{Gimenez, P.; Sengupta, I.; Srinivasan, H.} {Minimal free resolutions for certain affine monomial curves}. {\em Contemporary Mathematics} \textbf{555} (2011) 87-95.

\bibitem{Herzog70}
{Herzog, J.} {Generators and relations of abelian semigroups and semigroup rings}. {\em Manuscripta Math.} 3 (1970), 175--193.

\bibitem{OjKa}
{Katsabekis, A., Ojeda, I.} {An indispensable classification of monomial curves in $\mathbb{A}^4(\Bbbk)$}. {\em Pacific J. Math.} 268 (2014), 96--116. 

\bibitem{Ka17}
{Katsabekis, A.} {Arithmetical rank of binomial ideals}. {\em Arch. Math.} \textbf{109} (2017), 323--334.

\bibitem{KT10}
{Katsabekis, A., Thoma, A.} \emph{Specializations of multigradings and
the arithmetical rank of lattice ideals}. {\em Comm. Algebra} \textbf{38} (2010), 1904--1918.

\bibitem{MO18}
{Matusevich, {L.F.}, Ojeda, I.} {Binomial ideals and congruences on $\mathbb{N}^n$}. In: Greuel, G.M., et al. (eds.) {\em Singularities, Algebraic Geometry, Commutative Algebra, and Related Topics Festschrift for Antonio Campillo on the Occasion of his 65th Birthday}, pp. 429--454. Springer, Berlin (2018)

\bibitem{MS05}
{Miller, E.; Sturmfels, B.} {\em Combinatorial Commutative Algebra}. Vol. 227 of Graduate Texts in Mathematics. Springer, New York. 2005.

\bibitem{OV2}
{Ojeda, I; Vigner\'on-Tenorio, A.} {Indispensable binomials in semigroups ideals.} {\em Proc. Amer. Math. Soc.} \textbf{138} (2010), 4205--4216.

\bibitem{RBT16} 
{{Rosales, J.C.}; {Branco, M.B.}; {Torr\~ao, D.}}
{The Frobenius problem for repunit numerical semigroups}, {\em Ramanujan J.} \textbf{40} (2016), 323--334.

\bibitem{RBT17} 
{{Rosales, J.C.}; {Branco, M.B.}; {Torr\~ao, D.}} {The Frobenius problem for Mersenne numerical semigroups}, {\em Math. Z.} \textbf{286} (2017), 741--749.

\bibitem{libro}  
{{Rosales, J.C.}; {Garc\'{\i}a-S\'{a}nchez, P.A.}} {\em Numerical semigroups}. Developments in Mathematics, vol.\textbf{20}, Springer, New York, (2009).

\bibitem{Stu96}
{Sturmfels, B.} {\em Gröbner bases and convex polytopes}, AMS University Lecture Series Vol. \textbf{8}, American Mathematical Society, Providence, RI, 1996.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsfonts,amsmath,amssymb,amsthm}
\usepackage[right=3cm,left=3cm,top=3cm,bottom=3cm]{geometry}
\usepackage{enumerate}
\usepackage{mathdots}
\usepackage{url}
\usepackage{xcolor}

\newtheorem{Theorem}{Theorem}
\newtheorem{Lemma}[Theorem]{Lemma}
\newtheorem{Proposition}[Theorem]{Proposition}
\newtheorem{Corollary}[Theorem]{Corollary}

\theoremstyle{definition}

\newtheorem{Example}[Theorem]{Example}

\newtheorem{Remark}[Theorem]{Remark}

\title[Minimal systems of generators for the ideals of monomial curves]{Minimal binomial systems of generators for the ideals of certain monomial curves}

\author{Manuel B. Branco}
\address{Departamento de Matem\'aticas, Universidade de \'Evora, 7000-671 \'Evora, Portugal}
\email{mbb@uevora.pt}
\author{Isabel Colaço}
\address{Departamento de Matem\'atica e Ci\^encias F\'{\i}sicas, Instituto Polit\'ecnico de Beja, 7800-295 Beja, Portugal}
\email{isabel.colaco@ipbeja.pt}
\author{Ignacio Ojeda}
\address{Departamento de Matem\'aticas, Universidad de Extremadura, 06071 Badajoz, Spain}
\email{ojedamc@unex.es}
\thanks{This reserach was partially supported by the Ministerio de Economía y Competitividad
(Spain) /FEDER-UE under grants PGC2018-096446-B-C21 and MTM2017-84890-P, by the Junta de Extremadura(Spain)/FEDER funds, research group FQM-024,}
\subjclass[2010]{Primary: 13P10, 20M14, secondary: 52B20}
\keywords{Binomial ideal, semigroup ideal, minimal system of generators, determinantal ideal, Gr\"obner basis, indispensability.}

\begin{document}

\begin{abstract}
Let $a, b$ and $n > 1$ be three positive integers such that $a$ and $\sum_{j=0}^{n-1} b^j$ are relatively prime. In this paper, we prove that the toric ideal $I$ associated to the submonoid of $\mathbb{N}$ generated by $\{\sum_{j=0}^{n-1} b^j\} \cup \{\sum_{j=0}^{n-1} b^j + a\, \sum_{j=0}^{i-2} b^j \mid i = 2, \ldots, n\}$ is determinantal. Moreover, we prove that for $n > 3$, the ideal $I$ has a unique minimal system of generators if and only if  $a < b-1$.
\end{abstract}

\maketitle

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Introduction}

Let $\Bbbk$ be a field and let $\mathcal{A} = \{a_1, \ldots, a_n\}$ be a set of positive integers. It is well-known that the kernel of the $\Bbbk-$algebra homomorphism \begin{equation}\label{ecu2} \varphi_{\mathcal{A}} : \Bbbk[x_1, \ldots, x_n] \to \Bbbk[t^{a_1}, \ldots, t^{a_n}];\ x_i \mapsto t^{a_i},\ i = 1, \ldots, n,\end{equation} where $x_1, \ldots, x_n$ and $t$ are indeterminates, is a binomial ideal (see	 \cite{Herzog70} or \cite{MO18} for a more recent reference). Clearly, $\ker(\varphi_\mathcal{A})$ is the defining ideal of a monomial curve.

Let $b$ be a positive integer and set $r_b(\ell)$ for the $\ell-$th repunit number in base $b$, that is, \[r_b(\ell) = \sum_{j=0}^{\ell-1} b^j.\] By convention, $r_b(0) = 0$. 

The main result in this paper is the explicit determination of a minimal system of binomial generators of $I := \ker \varphi_{\mathcal{A}}$ for \[\mathcal{A} = \{a_i := r_b(n) + a\, r_b(i-1)\ \mid\ i = 1, \ldots, n\},\] where $a$ and $n > 1$ are positive integers. We prove that $I$ is minimally generated by the $2 \times 2$ minors of the matrix \begin{equation}\label{ecu1}
X:= \left(
\begin{array}{cccc} 
x_1^b  & \ldots & x_{n-1}^b & x_n^b \smallskip \\ x_2 & \ldots & x_n & x_1^{a+1}
\end{array}
\right), 
\end{equation}
provided that $\gcd(a_1, \ldots, a_n) = \gcd(a, r_b(n))$ is equal to $1$. In this case, as an immediate consequence, we have that the so-called binomial arithmetical rank of $I$ (see, e.g. \cite{Ka17}) is equal to $\binom{n}{2}$. 

Furthermore, we obtain that $2 \times 2-$minors of $X$ forms a minimal Gr\"obner basis with respect to a family of $\mathcal{A}-$graded reverse lexicographical term orders on $\Bbbk[x_1, \ldots, x_n]$ (Theorem \ref{Prop GB2}) and, applying \cite[Corollary 14]{OV2}, we conclude that for $n > 3$, the ideal $I$ has a unique minimal system of generators if and only if and $a < b-1$ (Corollary \ref{Cor GB2}).

The submonoids of $\mathbb{N}$ generated by $\mathcal{A}$ are studied in detail in \cite{BCO1} as a generalization of the numerical semigroups introduced by D. Torr\~ao et al. in \cite{RBT16} and \cite{RBT17}; in this context, Corollary \ref{Cor GB2} provides a minimal presentation of the submonoid of $\mathbb{N}$ generated by $a_1, \ldots, a_n$, providing an original result not considered in Torr\~ao's PhD thesis.

To achieve our main result (Theorem \ref{Prop GB2}), we first compute the ideal $J$ of the projective monomial curve defined by the kernel of the $\Bbbk-$algebra homomorphism \begin{equation}\label{ecu3}\Bbbk[x_1, \ldots, x_n] \to \Bbbk[t^{r_b(0)}\, s, \ldots, t^{r_b(n-1)}\, s];\quad x_i \mapsto t^{r_b(i-1)}\, s,\ i = 1, \ldots, n,\end{equation} where $s$ is also an indeterminate. This intermediate result (Proposition \ref{Prop GB1}) has its own interest, since it exhibits another family of semigroup ideals that are determinantal and have unique minimal system of binomial generators (Corollary \ref{Cor GB1}).

Throughout the paper, we keep the notation established in this introduction. Moreover, since the case $ n = 2 $ is trivial and the case $ n = 3 $ is well known for any $a_1, a_2$ and $a_3$ (see \cite{Herzog70}), we suppose that $n > 3$ whenever necessary.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Gr\"obner bases and minimal generators for $J$}\label{S2}

In the following, we write $a_{n+k} := r_b(n) + a\, r_b(n+k-1)$ for every $k \geq 1$. Observe that $a_{n+1} = (1+a)\, r_b(n)$.

\begin{Lemma}\label{lema2}
For each pair of positive integers $j$ and $k$, it holds that \[b\, a_j + a_{j+k} = b\, a_{j+k-1} + a_{j+1}.\]
\end{Lemma}

\begin{proof}
Since $a_{j+k} = a_1 + a\, r_b(j+k-1) = a_1 + a\, (r_b(j-1) +  b^{j-1}\, r_b(k)) = a_j + a\, b^{j-1}\, r_b(k)$, we conclude that 
\begin{align*}
b\, a_j + a_{j+k} & = b\, a_j + a_j + a\, b^{j-1}\, r_b(k)= \\ & = b\, a_j + a_j + a\, b^ {j-1}\, (b\, r_b(k-1)+1) = \\ & = b\, (a_j + a\, b^{j-1}\, r_b(k-1)) + a_j + a\, b^{j-1} = \\ & =  b\, a_{j+k-1} + (a_1 + a\, r_b(j-1)) + a\, b^{j-1} = \\ & =  b\, a_{j+k-1} + a_{j+1},
\end{align*}
as claimed.
\end{proof}

Let $\prec_i$ be the term order on $\Bbbk[x_1, \ldots, x_n]$ defined by the following matrix
\[
M:=\left(
\begin{array}{ccc|c|ccc}
a_1 & \ldots & a_i & a_{i+1} & a_{i+2} & \ldots & a_n\\ \cline{1-7}
0 &  & -1 & 0 & 0 & \ldots & 0 \\ 
& \iddots &  & \vdots & \vdots & & \vdots \\ 
-1 &  & 0 & 0 & 0 & \ldots & 0 \\ \cline{1-7}
0 & \ldots &  0 & 0 & 0 &  & -1 \\ 
\vdots & & \vdots & \vdots & & \iddots & \\ 
0 & & 0 & 0 & -1 & & 0
\end{array}
\right).
\]
We observe that $\prec_i$ is the $\mathcal{A}-$graded reverse lexicographical term order on $\Bbbk[x_1, \ldots, x_n]$ induced by $x_i \prec_i x_{i-1} \prec_i \ldots \prec_i x_1 \prec_i x_n \prec_i \ldots \prec_i x_{i+1};$ in particular, $x_i$ is the smallest variable for $\prec_i$.

\begin{Lemma}\label{lema3}
If $j \in \{1, \ldots, n-2\}$ and $k \in \{j+1, \ldots, n-1\}$, then \[x_j^b x_{k+1} \prec_i x_{j+1} x_k^b\] if and only if  $i \leq j$ or $k+1 \leq i$.
\end{Lemma}

\begin{proof}
By Lemma \ref{lema2}, $b\, a_j + a_{k+1} = a_{j+1} + b\, a_k$, so we just need to decide what the variable $x_j, x_{j+1}, x_k$ or $x_{k+1}$ is cheapest for the order defined by the last $n-1$ rows of $M$. Since $j < j+1 \leq k < k+1$, according to the definition of $\prec_i$, the variable $x_{k+1}$ is cheaper than the other three when $j \leq i$ or $k+1 \leq i$; thus, $x_j^b x_{k+1} \prec_i x_{j+1} x_k^b$ in these cases. Conversely, if $j+1 \leq i \leq k$, then either $x_k$ or $x_{j+1}$ is cheaper than the others if $k=i$ or $k \neq i$, respectively. Therefore $x_j^b x_{k+1} \succ_i x_{j+1} x_k^b$ when $j+1 \leq i \leq k$, and we are done.
\end{proof}

Let $I_2(Y)$ be the ideal of $\Bbbk[x_1, \ldots, x_n]$ generated by the $2 \times 2-$minors of \[
Y:= \left(
\begin{array}{ccccc} 
x_1^b & x_2^b & \ldots & x_{n-1}^b \smallskip \\ x_2 & x_3 & \ldots & x_n
\end{array}
\right).
\]
Let $\mathcal{G}^{(i)}_1, \mathcal{G}^{(i)}_2$ and $\mathcal{G}^{(i)}_3$ be defined as follows
\[
\mathcal{G}^{(i)}_1 = \left\{ \underline{x_{j+1} x_k^b} - x_j^b x_{k+1} \ \mid\ j \in \{i, \ldots, n-2\},\ k \in \{j+1,\ldots, n-1\}\right\},
\]
\[
\mathcal{G}^{(i)}_2 = \left\{ \underline{x_{j+1} x_k^b} - x_j^b x_{k+1} \ \mid\ j \in \{1, \ldots, i-2\},\ k \in \{j+1,\ldots, i-1\}\right\},
\]
\[
\mathcal{G}^{(i)}_3 = \left\{ \underline{x_j^b x_{k+1}} - x_{j+1} x_k^b \ \mid\ j \in \{1, \ldots, i-1\},\ k \in \{i,\ldots, n-1\}\right\}
\]
and let $\mathcal{G}_Y^{(i)}$ be equal to $\mathcal{G}^{(i)}_1 \cup \mathcal{G}^{(i)}_2 \cup \mathcal{G}^{(i)}_3$.

Notice that, by Lemma \ref{lema3}, the underlined monomials are the leading terms with respect to $\prec_i$ of the corresponding binomials.

\begin{Proposition}\label{Prop GB1}
With the above notation, the set $\mathcal{G}_Y^{(i)}$ is the reduced Gr\"obner basis of $I_2(Y)$ with respect to $\prec_i$. In particular, the cardinality of $\mathcal{G}_Y^{(i)}$ is $\binom{n-1}2$.
\end{Proposition}

\begin{proof}
First, let us see that $\mathcal{G}_Y^{(i)}$ is a Gr\"obner basis. By the Buchberger’s Criterion (see, e.g. \cite[Theorem 3.3]{MS05}), it suffices to verify that each S-pair of elements in $\mathcal{G}_Y^{(i)}$ can be reduced to zero by $\mathcal{G}_Y^{(i)}$ using the division algorithm. To do this we distinguish several cases:
\begin{itemize}
\item Let $f \in \mathcal{G}^{(i)}_1$, that is to say, $f = x_{j+1} x_k^b - x_j^b x_{k+1}$, for some $j \in \{i, \ldots, n-2\}$ and $k \in \{j+1,\ldots, n-1\}.$
\begin{itemize}
\item[$\circ$] Let $g = x_{l+1} x_m^b - x_l^b x_{m+1} \in \mathcal{G}^{(i)}_1$. If $\gcd(x_{j+1} x_k^b, x_{l+1} x_m^b) = 1$, then $S(f,g)$ reduces to zero with respect to $\{f,g\} \subset \mathcal{G}_Y^{(i)}$. Otherwise, $j = l, j+1 = m, k = l+1$ or $k = m$. If \framebox{$j=l$} then $S(f,g) = x_m^b (- x_j^b x_{k+1}) - x_k^b (- x_j^b x_{m+1}) = x_j^b (x_k^b x_{m+1} - x_m^b x_{k+1})$ reduces to zero with respect to $\mathcal{G}_Y^{(i)}$. If \framebox{$j+1 = m$}, then \[S(f,g) = x_{l+1} x_{j+1}^{b-1}(- x_j^b x_{k+1}) - x_k^b(-x_l^b x_{j+2}).\] Now, since $i \leq j < j+1 \leq k < k + 1$ and $i \leq l < l+1 \leq m=j+1 < j+2$, the leading term of $S(f,g)$ with respect to $\prec_i$ is $x_k^b x_l^b x_{j+2}$. Then $S(f,g) = x_l^b ( \underline{x_k^b x_{j+2}} - x_{j+1}^b x_{k+1}) + x_{j+1}^{b-1} x_{k+1}  (x_l^b x_{j+1} - x_{l+1} x_j^b)$ reduces to zero with respect to $\mathcal{G}_Y^{(i)}$. By symmetry, the case \framebox{$k=l+1$} is completely similar to the latter one.  Finally, if \framebox{$k=m$}, then $S(f,g) = -x_{l+1}(-x_j^b x_{k+1}) - x_{j+1}(-x_l^b x_{k+1}) =  x_{k+1}(x_j^b x_{l+1} - x_{j+1} x_l^b)$ reduces to zero with respect to $\mathcal{G}_Y^{(i)}$.
\item[$\circ$] Let $g = x_{l+1} x_m^b - x_l^b x_{m+1} \in \mathcal{G}^{(i)}_2$. If $\gcd(x_{j+1} x_k^b, x_{l+1} x_m^b) = 1$, then $S(f,g)$ reduces to zero with respect to $\{f,g\} \subset \mathcal{G}_Y^{(i)}$. Otherwise, $j = l, j+1 = m, k = l+1$ or $k = m$. First of all, we observe that the cases $j=l$ and $k=m$ produce the same S-polynomial as in the corresponding case for $g \in \mathcal{G}^{(i)}_1$; so, we just focus on the cases $j+1 =m$ and $k=l+1$. If \framebox{$j+1 =m$}, then $i \leq j < j+1 \leq k < k + 1$ and $l < l+1 \leq m=j+1 < j+2 = m-1 \leq i$, therefore $i < j+1 = m < i$, a contradiction. Finally, if \framebox{$k=l+1$}, then $i \leq j < j+1 \leq k < k + 1$ and $l < l+1=k \leq m < m+1 \leq i$, so $i < k = l+1 < i$, a contradiction again.
\item[$\circ$] Let $g = x_l^b x_{m+1} - x_{l+1} x_m^b \in \mathcal{G}^{(i)}_3$. If $\gcd(x_{j+1} x_k^b, x_l^b x_{m+1}) = 1$, then $S(f,g)$ reduces to zero with respect to $\{f,g\} \subset \mathcal{G}_Y^{(i)}$. Otherwise, $j+1 = l, j = m, k = l$ or $k = m+1$. If \framebox{$j+1 = l$}, then $i \leq j < j+1 = l \leq k < k + 1$ and $l \leq i-1$; so $i < j+1 = l \leq i-1$, a contradiction. If \framebox{$j = m$} (or \framebox{$k = l$}, respectively) then $S(f,g) = x_m^b(x_k^b x_{l+1} - x_l^b x_{k+1})$ (or $S(f,g) = x_{k+1}(x_{j+1} x_m^b - x_j^b x_{m+1}),$  respectively) reduces to zero with respect to $\mathcal{G}_Y^{(i)}$. Finally, if \framebox{$k=m+1$}, then \[S(f,g) = x_l^b(- x_j^b x_{m+2}) - x_{j+1} x_{m+1}^{b-1} (- x_{l+1} x_{m}^b).\] Now, since $i \leq j < j+1 \leq k = m+1 < k + 1$, $l \leq i-1$ and $i \leq m$, then $x_{l+1}$ or $x_{m-1}$ is cheaper than the others for the order induced by the last $n-1$ rows of the matrix $M$, therefore leading term of $S(f,g)$ is $x_l^b x_j^b x_{k+1}$ and thus, $S(f,g) = - x_j^b (\underline{x_l^b x_{m+2}} - x_{l+1} x_{m+1}^b) - x_{l+1} x_{m+1}^{b-1}(x_j^b x_{m+1}- x_{j+1} x_{m}^b)$ reduces to zero with respect to $\mathcal{G}_Y^{(i)}$.
\end{itemize}
\item Let $f \in \mathcal{G}^{(i)}_2$, that is to say, $f = x_{j+1} x_k^b - x_j^b x_{k+1}$, for some $j \in \{1, \ldots, i-2\}$ and $k \in \{j+1,\ldots, i-1\}.$
\begin{itemize}
\item[$\circ$] Let $g = x_{l+1} x_m^b - x_l^b x_{m+1} \in \mathcal{G}^{(i)}_2$. If $\gcd(x_{j+1} x_k^b, x_{l+1} x_m^b) = 1$, then $S(f,g)$ reduces to zero with respect to $\{f,g\} \subset \mathcal{G}_Y^{(i)}$. Otherwise, $j = l, j+1 = m, k = l+1$ or $k = m$. If \framebox{$j=l$} (or \framebox{$k = m$}, respectively), then $S(f,g) = x_j^b(x_k^bx_{m+1}-x_{k+1}x_m^b)$ (or $S(f,g) = x_{k+1}(x_{l+1}  x_j^b- x_l^bx_{j+1})$, respectively) reduces to zero with respect to $\mathcal{G}_Y^{(i)}$. If \framebox{$j+1 = m$}, then \[S(f,g) = x_{l+1} x_m^{b-1}(- x_{m-1}^b x_{k+1}) - x_k^b (- x_l^b x_{m+1})\] and, since $l+1 \leq m=j+1 \leq k \leq i-1$, the leading term of $S(f,g)$ is $ x_{l+1} x_m^{b-1} x_{m-1}^b x_{k+1}$. Thus, $S(f,g) = -x_m^{b-1} x_{k+1} (\underline{x_{l+1} x_{m-1}^b} - x_l^b x_m) + x_l^b(x_k^b x_{m+1} - x_m^{b} x_{k+1})$ reduces to zero with respect to $\mathcal{G}_Y^{(i)}$; observe that $l < l+1 \leq m$ implies that the leading term of $x_{l+1} x_{m-1}^b - x_l^b x_m$ is actually $x_{l+1} x_{m-1}^b$. Finally, by symmetry, the case \framebox{$k=l+1$} is completely similar to the latter one.
\item[$\circ$] Let $g = x_l^b x_{m+1} - x_{l+1} x_m^b \in \mathcal{G}^{(i)}_3$. If $\gcd(x_{j+1} x_k^b, x_l^b x_{m+1}) = 1$, then $S(f,g)$ reduces to zero with respect to $\{f,g\} \subset \mathcal{G}_Y^{(i)}$. Otherwise, $j+1 = l, j = m, k = l$ or $k = m+1$. If \framebox{$j+1=l$}, then \[S(f,g) = x_l^{b-1} x_{m+1} (-x_{l-1}^b x_{k+1}) -x_k^b ( - x_{l+1} x_m^b).\] Furthermore, since $l = j+1 \leq k \leq i-1$ and $l \leq i-1 < i \leq m < m+1$, we have that the leading term is $x_l^{b-1} x_{m+1} x_{l-1}^b x_{k+1}$ if $k=l$ and $x_k^b x_{l+1} x_m^b$ otherwise. In ther first case, $S(f,g) =  x_{m+1}x_{k-1}^b +x_k x_m^b$ reduces to zero with respecto to $\mathcal{G}_Y^{(i)}$. In the second case, $S(f,g) = x_m^b (\underline{x_k^b x_{l+1}} - x_{k+1} x_l^b) +  x_l^{b-1}  x_{k+1} (x_m^b x_l -   x_{m+1} x_{l-1}^b)$ reduces to zero with respect to $\mathcal{G}_Y^{(i)}$. If \framebox{$j = m$} (or \framebox{$k = l$}, respectively) then $S(f,g) = x_m^b(x_k^b x_{l+1} - x_l^b x_{k+1})$ (or $S(f,g) = x_{k+1}(x_{j+1} x_m^b - x_j^b x_{m+1}),$  respectively) reduces to zero with respect to $\mathcal{G}_Y^{(i)}$. Finally, if \framebox{$k=m+1$}, then 
$j+1 \leq k = m+1 \leq i-1, l \leq i-1$ and $i \leq m$; so, $m+1 < i \leq m$, a contradiction.
\end{itemize}
\item Let $f \in \mathcal{G}^{(i)}_3$, that is to say, $f = x_j^b x_{k+1} - x_{j+1} x_k^b,$ for some $j \in \{1, \ldots, i-1\}$ and $k \in \{i,\ldots, n-1\}$.
\begin{itemize}
\item[$\circ$] Let $g = x_l^b x_{m+1} - x_{l+1} x_m^b \in \mathcal{G}^{(i)}_3$. If $\gcd(x_j^b x_{k+1}, x_l^b x_{m+1}) = 1$, then $S(f,g)$ reduces to zero with respect to $\{f,g\} \subset \mathcal{G}_Y^{(i)}$. Otherwise, $j = l, j = m+1, k+1 = l$ or $k = m$. Since $j \leq i-1 < i \leq k$ and $l \leq i-1 < i \leq m$, the cases \framebox{$j = m+1$} and \framebox{$k+1 = m$} cannot occur. If \framebox{$j = l$} (or \framebox{$k=m$}, respectively), then $S(f,g) = x_{l+1}(x_{k+1} x_m^b - x_{m+1}x_k^b)$ (or $S(f,g) = x_m^b (x_l^b x_{j+1} - x_j^b x_{l+1} )$, respectively) reduces to zero with respect $\mathcal{G}_Y^{(i)}$.
\end{itemize}
\end{itemize}
Once we know that $\mathcal{G}_Y^{(i)}$ is Gr\"obner basis, it is immediate to see that it is reduced since the leading term of $f \in \mathcal{G}_Y^{(i)}$ does not divide any other monomial that appears in a binomial of $\mathcal{G}_Y^{(i)} \setminus \{f\}$.

It remains to prove that $\mathcal{G}_Y^{(i)}$ generates $I_2(Y)$. Clearly, $\mathcal{G}_Y^{(i)}$ is contained in the set of $2 \times 2-$minors of $Y$. Moreover, since the cardinality of $\mathcal G_1, \mathcal G_2$ and $\mathcal G_3$ are 
\[\Big((n-1)-i\Big) + \Big((n-1)-i-1\Big) + \ldots + 1 = \binom{n-i}2,\]
\[\Big((i-1)-1\Big) + \Big((i-1)-2 \Big) + \ldots + 1 = \binom{i-1}2\]
and
\[(i-1)(n-i),\]
respectively, we have that the cardinality of $\mathcal{G}_Y^{(i)}$ is equal to $\binom{n-1}2$ which is the number of $2 \times 2-$minors of $Y$. Therefore $\mathcal{G}_Y^{(i)}$ generates $I_2(Y)$ and we are done.
\end{proof}

\begin{Example}\label{Ex5}
We observe that the reduced Gr\"obner basis, $\mathcal{G}_Y^{(i)}$, of $I_2(Y)$ with respect to $\prec_i$ is not an universal Gr\"obner basis. For example, if $n=b=5$ and $\prec$ is the term order defined by 
\[
\left(
\begin{array}{ccccc}
a_1 & a_2 & a_3 & a_4 & a_5 \\
0 & 0 & -1 & 0 & 0 \\
0 & 0 & 0 & 0 & -1 \\
0 & 0 & 0 & -1 & 0 \\
0 & -1 & 0 & 0 & 0
\end{array}
\right),
\]
then one can check (using, for example, Singular \cite{singular}) that the reduced Gr\"obner basis of the ideal $I_2(Y)$ with respect to $\prec$ has eight generators; but $\mathcal{G}_Y^{(i)}$ contains $\binom{5-1}2 = 6$ binomials only. 

Alternatively, one can see that $\mathcal{G}_Y^{(i)}$ is not an universal Gr\"obner basis of $I_2(Y)$ by using \cite[Theorem 4.1]{BR15}.
\end{Example}

We now consider the $2 \times n-$integer matrix $B$ whose $j-$th column is \[\mathbf{a}_j := \left(\begin{array}{c} r_b(j-1)\\ 1 \end{array}\right),\quad j = 1, \ldots, n.\] 

\begin{Remark}
Observe that $a_j = \big(a, r_b(n)\big) \cdot \mathbf{a}_j,$ for every $j = 1, \ldots, n.$
\end{Remark}

Notice that the semigroup ideal associated to $\{\mathbf{a}_1, \ldots, \mathbf{a}_n\}$ is equal to $J$; indeed, $J$ is the kernel of \eqref{ecu3}.

\begin{Corollary}\label{Cor GB1}
The ideal $J$ is minimally generated by the $2 \times 2-$minors of $Y$. Moreover, $J$ has a unique minimal system of binomial generators.
\end{Corollary}

\begin{proof}
Let $I_2(Y)$ the ideal generated by the $2 \times 2-$minors of $Y$. Since $b\, \mathbf{a}_j + \mathbf{a}_{k+1} = \mathbf{a}_{j+1} + b\, \mathbf{a}_k$ for every $j$ and $k$, we have that $I_2(Y) \subseteq J$. 

Conversely, let $C$ be the $(n-2) \times n-$matrix 
\[
\left(
\begin{array}{ccccccccccc}
b & -1 & -b &  1 & 0 & \ldots & 0 & 0 & 0 & 0 \\
0 &  b & -1 & -b & 1 & \ldots & 0 & 0 & 0 & 0 \\
0 & 0 &  b & -1 & -b & \ldots & 0 & 0 & 0 & 0 \\
0 & 0 & 0 &  b & -1 & \ldots & 0 & 0 & 0 & 0 \\
0 & 0 & 0 & 0 & b  & \ldots & 0 & 0 & 0 & 0 \\
\vdots & \vdots & \vdots & \vdots & \vdots & \ddots & \vdots & \vdots & \vdots & \vdots \\
0 & 0 & 0 & 0 & 0 & \ldots & b & -1 & -b & 1\\
0 & 0 & 0 & 0 & 0 & \ldots & 0 & b & -(b+1) & 1\\
\end{array}
\right)
\]
and let $I_{C}$ be the ideal of $\Bbbk[x_1, \ldots, x_n]$ generated by \[\{ \mathbf{x}^{\mathbf{u}_+} - \mathbf{x}^{\mathbf{u}_-}\ \mid\ \mathbf{u}\ \text{is a row of}\ {C} \},\] where $\mathbf{u}_+$ and $\mathbf{u}_-$ denote the positive and negative parts of $\mathbf{u}$, respectively. Clearly, $I_{C} \subseteq I_2(Y)$.

Now, since the determinant of the submatrix of $C$ consisting in the last $n-2$ columns is $1$, the rows of $C$ generates a rank $n-2$ subgroup $G_{C}$ of $\mathbb{Z}^n$ such that $\mathbb{Z}^n/G_{C}$ is torsion free. Moreover, since $B\, C^\top = 0$, we conclude that the rows of $C$ generate $\ker_\mathbb{Z}(B)$. Therefore, by \cite[Lemma 7.6]{MS05}, \[J = I_{C}: \Big(\prod_j x_j \Big)^\infty  \subseteq I_2(Y) : \Big(\prod_j x_j \Big)^\infty .\] By Proposition \ref{Prop GB1} and \cite[Theorem 3.1]{BSR99}, we have that $I_2(Y):x_i^\infty = I_2(Y)$ for every $i = 1, \ldots, n$. So, $I_2(Y) : (\prod_j x_j )^\infty = I_2(Y)$ and, consequently, $J \subseteq  I_2(Y)$ as desired. 

Finally, by Proposition \ref{Prop GB1}, we conclude that the $2 \times 2-$minors of $Y$ form a minimal system generators of $J$ and,  \cite[Corollary 14]{OV2}, we conclude that $J$ has a unique minimal system of binomial generators.
\end{proof}

We recall that semigroup ideals minimally generated by a Graver basis have unique minimal system of binomials generators (see \cite[Corollary 16]{OV2}). As Graver bases are in particular universal Gr\"obner bases (see \cite[Proposition 4.11]{Stu96}), by Example \ref{Ex5}, we can assure the minimal system of binomial generators of $J$ is not a Graver basis. %Therefore, we exhibit a new family of semigroup ideal ideals having unique minimal system of binomial generators.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Gr\"obner basis and minimal generators for $I$}

We keep the notation from the Introduction and Section \ref{S2} and we set
$\mathcal G_4$ to be equal to \[\left\{\underline{x_1^{a+1} x_l^b} - x_{l+1} x_n^b \ \mid\ l = 1, \ldots, i-1  \right\}\bigcup \left\{\underline{x_{l+1} x_n^b} - x_1^{a+1} x_l^b \mid\ l = i, \ldots, n-1  \right\},\] where the underlined monomials again highlight the leading terms with respect to $\prec_i$ of the corresponding binomials.

Let $I_2(X)$ be the ideal of $\Bbbk[x_1, \ldots, x_n]$ generated by the $2 \times 2-$minors of the matrix $X$ defined in \eqref{ecu1}.

\begin{Theorem}\label{Prop GB2}
The set $\mathcal{G} = \mathcal{G}_Y^{(i)} \cup \mathcal{G}^{(i)}_4$ is a minimal Gr\"obner basis of $I_2(X)$ with respect to $\prec_i$. In particular, the cardinality of $\mathcal G$ is $\binom{n}2$.
\end{Theorem}

\begin{proof}
Proceeding as in the proof of Proposition \ref{Prop GB1}, we first need to prove that $S(f,g)$ reduces to zero with respect to $\mathcal{G}^{(i)}$, for every $f, g \in \mathcal{G}^{(i)}$. However, since, by Proposition \ref{Prop GB1}, $\mathcal{G}_Y^{(i)}$ is already a Gr\"obner basis with respect to $\prec_i$ and the leading terms with respect to $\prec_i$ of the binomials in $\mathcal{G}^{(i)}_4$ are relatively prime, it suffices to prove that $S(f,g)$ reduces to zero with respect to $\mathcal{G}^{(i)}$, for every $f \in \mathcal{G}_Y^{(i)}$ and $g \in \mathcal{G}^{(i)}_4$. To do this we distinguish three cases:
\begin{itemize}
\item $f \in \mathcal{G}^{(i)}_1 = \big\{ \underline{x_{j+1} x_k^b} - x_j^b x_{k+1} \ \mid\ j \in \{i, \ldots, n-2\},\ k \in \{j+1,\ldots, n-1\}\big\}.$ If $j \neq l$ and $k \neq l+1$, then the leading terms of $f$ and $g$ are relatively prime and there is nothing to prove. So, it suffices to consider the cases $j = l$ or 
$k = l+1$.
\begin{itemize}
\item[$\circ$] If $j = l$, then $l \geq i$, otherwise the leading terms of $f$ and $g$ are relatively prime, and $S(f,g) = x_n^b (-x_l^b x_{k+1}) - x_k^b (-x_1^{a+1} x_l^b) = -x_l^b( x_{k+1} x_n - x_1^{a+1} x_l^b)$ reduces to zero with respect to $\mathcal{G}^{(i)}_4$. 
\item[$\circ$] If $k = l+1$ then $n-2 \geq k-1 = l \geq j \geq i$, otherwise the leading terms of $f$ and $g$ are relatively prime, and $S(f,g) = x_n^b (-x_j^b x_{l+2}) - x_{j+1} x_{l+1}^{b-1} (-x_1^{a+1} x_l^b) = \underline{- x_j^b x_n^b x_{l+2}} + x_1^{a+1}  x_{j+1} x_{l+1}^{b-1} x_l^b$. Observe that the leading term of $S(f,g)$ is divisible by the leading term of $h := \underline{x_n^b x_{l+2}} - x_1^{a+1} x_{l+1}^b  \in \mathcal{G}^{(i)}_4$. So, $S(f,g) = x_j^b h - x_j^b x_1^{a+1} x_{l+1}^b + x_1^{a+1}  x_{j+1} x_{l+1}^{b-1} x_l^b = x_j^b h - x_1^{a+1} x_{l+1}^{b-1} (x_j^b x_{l+1} - x_{j+1}x_l^b)$. Now, since $x_j^b x_{l+1} - x_{j+1}x_l^b \in \mathcal{G}_Y^{(i)}$, we are done. 
\end{itemize}
\item $f \in \mathcal{G}^{(i)}_2 = \big\{ \underline{x_{j+1} x_k^b} - x_j^b x_{k+1} \ \mid\ j \in \{1, \ldots, i-2\},\ k \in \{j+1,\ldots, i-1\}\big\}$. If $j+1 \neq l$ and $k \neq l$, then the leading terms of $f$ and $g$ are relatively prime and there is nothing to prove. So, it suffices to consider the cases $j = l -1$ or $k = l$.
\begin{itemize}
\item[$\circ$] If $j+1 = l$, then $1 \leq j = l-1 < k \leq i-1$, otherwise the leading terms of $f$ and $g$ are relatively prime, and $S(f,g) = x_1^{a+1} x_l^{b-1} (-x_{l-1}^b x_{k+1}) - x_k^b  (- x_{l+1} x_n^b) = x_k^b x_{l+1} x_n^b - x_1^{a+1} x_{l-1}^b x_l^{b-1} x_{k+1}$. If $l=k$, then the S-polynomial $S(f,g) =  x_k^b x_{k+1} x_n^b - x_1^{a+1} x_{k-1}^b x_k^{b-1} x_{k+1} = - x_k^{b-1} x_{k+1}(\underline{x_1^{a+1} x_{k-1}^b} - x_k x_n^b)$ reduces to zero with respect to $\mathcal{G}^{(i)}_4$; otherwise, the leading term of $S(f,g)$ is $x_k^b x_{l+1} x_n^b$ which is divisible by the leading term of $h := \underline{x_k^b x_{l+1}} - x_{k+1} x_l^b \in \mathcal{G}^{(i)}_4$. 
So, $S(f,g) = x_n^b h + x_n^b x_{k+1} x_l^b  - x_1^{a+1} x_{l-1}^b x_l^{b-1} x_{k+1} = x_n^b h - x_l^{b-1} x_{k+1}(x_1^{a+1} x_{l-1}^b - x_n^b x_l).$ Now, since $x_1^{a+1} x_{l-1}^b - x_n^b x_l \in \mathcal{G}^{(i)}_4$, we are done.
\item[$\circ$] If $k=l$, then $1 \leq j < k = l \leq i-1$, otherwise the leading terms of $f$ and $g$ are relatively prime, and $S(f,g) = x_1^{a+1} (- x_j^b x_{l+1} ) - x_{j+1} (- x_{l+1} x_n^b )= - x_{l+1}( x_1^{a+1} x_j^b - x_{j+1} x_n^b).$ Now, since $x_1^{a+1} x_j^b - x_{j+1} x_n^b \in \mathcal{G}^{(i)}_4$, we are done.
\end{itemize}
\item $f \in \mathcal{G}^{(i)}_3 =  \big\{ \underline{x_j^b x_{k+1}} - x_{j+1} x_k^b \ \mid\ j \in \{1, \ldots, i-1\},\ k \in \{i,\ldots, n-1\}\big\}$. If $j \neq 1, j \neq l, k \neq l$ and $k \neq n-1$, then the leading terms of $f$ and $g$ are relatively prime and there is nothing to prove. So, is suffices so consider the cases $j=1, j = l, k=l$ or $k = n-1$.
\begin{itemize}
\item[$\circ$] If $j=1$, then, in particular, $l < i$, otherwise the leading terms of $f$ and $g$ are relatively prime. Now, if $a+1 \geq b$, then $S(f,g)= \underline{x_1^{a+1-b} x_l^b (- x_2 x_k^b)} - x_{k+1} x_{l+1} x_n^b$ and its leading term is divisible by the leading term of $h := \underline{x_l^b x_2} - x_{l+1} x_1^b \in \mathcal{G}^{(i)}_2 \cup \mathcal{G}^{(i)}_3$; then $S(f,g) = x_1^{a+1-b} x_k^b h - x_1^{a+1-b} x_k^b (- x_{l+1} x_1^b) - x_{k+1} x_{l+1} x_n^b = x_1^{a+1-b} x_k^b h - x_{l+1} (\underline{x_{k+1} x_n^b} - x_1^{a+1} x_k^b)$ which reduces to zero with respect to $\mathcal{G}^{(i)}_2 \cup \mathcal{G}^{(i)}_3 \cup \mathcal{G}^{(i)}_4$. Otherwise, if $a+1 < b$, then $S(f,g) = \underline{x_l^b (-x_2 x_k^b)} - x_1^{b-a-1} x_{k+1} (- x_{l+1} x_n^b)$ and its leading term is divisible by the leading term of $h := \underline{x_l^b x_2} - x_{l+1} x_1^b \in \mathcal{G}^{(i)}_2 \cup \mathcal{G}^{(i)}_3$; then $S(f,g) = x_k^b h - x_k^b (-x_{l+1} x_1^b) - x_1^{b-a-1} x_{k+1}(- x_{l+1} x_n^b) = x_k^b h - x_1^{b-a-1} x_{l+1} (x_{k+1} x_n^b-x_1^{a+1} x_k^b)$ which reduces to zero with respect to $\mathcal{G}^{(i)}_2 \cup \mathcal{G}^{(i)}_3 \cup \mathcal{G}^{(i)}_4$, too.
\item[$\circ$] If $j=l$, then $1 \leq l \leq i-1$, otherwise the leading terms of $f$ and $g$ are relatively prime, and $S(f,g)= x_1^{a+1} (- x_{l+1} x_k^b) - x_{k+1} (-x_{l+1} x_n^b) = -x_{l+1} (x_{k+1} x_n^b - x_1^{a+1} x_k^b)$ which reduces to zero with respect to $\mathcal{G}^{(i)}_4$.
\item[$\circ$] If $k=l$, then $i \leq l \leq n-1$, otherwise the leading terms of $f$ and $g$ are relatively prime, and $S(f,g) = x_n^b(-x_{j+1} x_l^b) - x_j^b (-x_1^{a+1} x_l^b) = x_l^b (\underline{x_1^{a+1} x_j^b} - x_{j+1} x_n^b)$ which reduces to zero with respect to $\mathcal{G}^{(i)}_4$.
\item[$\circ$] If $k=n-1$, then $1 \leq j < i \leq l < n$, otherwise the leading terms of $f$ and $g$ are relatively prime. In this case, $S(f,g)= x_{l+1} x_n^{b-1} (-x_{j+1} x_{n-1}^b) \underline{- x_j^b (-x_1^{a+1} x_l^b)}$ and, since the leading term of $S(f,g)$ is divisible by the leading term of $h:=\underline{x_1^{a+1} x_j^b} - x_{j+1} x_n^b \in \mathcal{G}^{(i)}_4$, we have that $S(f,g) = x_l^b h - x_l^b (- x_{j+1} x_n^b) +  x_{l+1} x_n^{b-1} (-x_{j+1} x_{n-1}^b) = x_l^b h - x_{j+1} x_n^{b-1}(\underline{x_{l+1} x_{n-1}^b} - x_l^b x_n)$, and since $\underline{x_{l+1} x_{n-1}^b} - x_l^b x_n$ belongs to $\mathcal{G}^{(i)}_1$, we are done.  
\end{itemize}
\end{itemize}
Now, since $S(f,g)$ reduces to zero with respect to $\mathcal{G}^{(i)}$ in all the three cases we conclude that $\mathcal{G}^{(i)}$ forms a Gr\"obner basis.

Once we know that $\mathcal{G}^{(i)}$ is a Gr\"obner basis, we observe that the leading terms of the binomials in $\mathcal{G}^{(i)}$ are not divisible by the leading term of any other binomial in $\mathcal{G}^{(i)}$ other than itself. That is to say, the Gr\"obner basis $\mathcal{G}^{(i)}$ is minimal.

Clearly, $\mathcal{G}^{(i)}$ is a subset of $2 \times 2-$minors of $X$. Moreover, its cardinality is equal to the cardinality of $\mathcal{G}_Y^{(i)},$ that is $\binom{n-1}2$, plus the cardinality, $n-1$, of $\mathcal{G}^{(i)}_4$. Therefore, $\mathcal{G}^{(i)}$ has cardinality equal to $\binom{n-1}2 + (n-1) = \binom{n}2$ which is the number of $2 \times 2-$minors of $X$. Hence we conclude that $\mathcal{G}^{(i)}$ generates $I_2(X)$ and we are done.
\end{proof}

\begin{Example}
The minimal Gr\"obner basis, $\mathcal{G}^{(i)}$, of $I_2(X)$ with respect to $\prec_i$ is not reduced in general. For example, if $n = 4, a = 3$ and $b = 3$, then one can see (using, e.g. Singular \cite{singular}) that the binomial $x_4^4 - x_1 x_2^4 x_3^2 $ belongs to the Gr\"obner basis of $I_2(X)$ with respect to $\prec_2$, however, $x_4^4 - x_1 x_2^4 x_3^2$ is not a minor of $X$.
\end{Example}

\begin{Corollary}\label{Cor GB2}
If $\gcd(a, r_b(n)) = 1$, then the ideal $I$ is minimally generated by the $2 \times 2-$minors of $X$. In this case, if $n > 3$, then $I$ has a unique minimal system of generators if and only if  and $a < b-1$.
\end{Corollary}

\begin{proof}
By Theorem \ref{Prop GB2}, to prove the first part of the statement it suffices to see that $I = I_2(X)$.

By Lemma \ref{lema2}, we have that $\varphi_\mathcal{A}(f) = 0$, for every $f \in \mathcal{G}^{(i)}$, where $\varphi_\mathcal{A}$ is the $\Bbbk-$algebra homomorphism define in \eqref{ecu2}. Therefore $I_2(X) \subseteq I$. Conversely, let $L$ be the $(n-1) \times n-$matrix 
\[\left(\begin{array}{cccccccc} b & -(b+1) & 1 & 0 & \ldots & 0 & 0 & 0 \\ 
0 & b & -(b+1) & 1 & \ldots & 0 & 0 & 0  \\ 
\vdots & \vdots & \vdots & \vdots &  & \vdots & \vdots & \vdots\\ 
0 & 0 & 0 & 0 & \ldots & b & -(b+1) & 1 \\
(a+1) & 0 & 0 & 0 & \ldots & 0 & b & -(b+1) 
\end{array}\right)\] and let $I_L$ be the ideal of $\Bbbk[x_1,  \ldots, x_n]$ generated by \[\left\{\mathbf{x}^{\mathbf{u}_+} - \mathbf{x}^{\mathbf{u}_-} \mid \mathbf{u}\ \text{is a row of}\ L\right\}.\] Clearly, $I_L \subseteq I_2(X)$.

On the one hand, a direct computation shows that the set of $(n-1) \times (n-1)-$minors of 
$L$ is equal to $\{a_1, \ldots, a_n\}$ and therefore, by \cite[Lemma 12.2]{Stu96}, \[ 
I_L : \big(\prod_{i=1}^n x_i \big)^\infty = I\] if and only if $\gcd(a_1, \ldots, a_n) = \gcd(a, r_b(n)) = 1$. On the other hand, by Theorem \ref{Prop GB2} and \cite[Theorem 3.1]
{BSR99}, we have that $I_2(X): \big(x_i^\infty\big) =  I_2(X)$ for every $i = 1, \ldots, 
n$, that is to say, $I_2(X): \big(\prod_{i=1}^n x_i\big)^\infty = I_2(X)$. Putting this 
together we conclude that \[ I_2(X) = I_2(X): \big(\prod_{i=1}^n x_i\big)^\infty \supseteq 
I_L : \big( \prod_{i=1}^n x_i\big)^\infty = I,\]
and thus $I = I_2(X)$ as claimed.

To prove the second part of the statement, we observe that, for every $i \neq n$, the non-leading term, $x_1^{a+1} x_l^b,$ of the binomial $\underline{x_{l+1} x_n^b} - x_1^{a+1} x_l^b  \in \mathcal{G}^{(i)}_4$ is divisible by the leading term of $\underline{x_1^b x_l} - x_2 x_{l-1}^b \in \mathcal{G}^{(i)}_3$, provided that $l \geq 3$ (otherwise no such binomial in $\mathcal{G}^{(i)}_3$ exists), if and only $a+1 \geq b$. Now, as these are the only divisibility relationships between the monomials of the binomials in $\mathcal{G}^{(i)}$, and $l \geq 3$ implicitly requires $n > 3$ , we obtain that for $n > 3$, $\mathcal{G}^{(i)}$ is reduced for every $\prec_i$ if and only $a < b - 1$, and, by \cite[Corollary 14]{OV2}, we conclude that for $n > 3,\ I$ has a unique minimal system of binomial generators if and only if and $a < b-1$.
\end{proof}

Notice that the condition $\gcd(a, r_b(n)) = 1$ cannot be avoided.

\begin{Example}
Let $n = 4, a = 3$ and $b = 2$. In this case, $a_1 = r_b(4) = 15, a_2 = 18, a_3 = 24$ and $a_4 = 36$. Clearly, $\gcd(a_1, a_2, a_3, a_4) = \gcd(a,r_b(4)) = 3$. By direct computation, one can check that $I$ is minimally generated by four binomials whereas $I_2(X)$ is minimally generated by six binomials. In particular, $I \neq I_2(X)$; in fact, one has that $I$ is a minimal prime of $I_2(X)$. 
\end{Example}

The following example shows the minimal system of generators of $I$ for $n=4$.

\begin{Example}
If $n=4$, then the ideal $I \subset \Bbbk[x_1,x_2,x_3,x_4]$ is minimally generated by 
\[x_2^{b+1} - x_1^b x_3, x_1^b x_4 - x_2 x_3^b, x_3^{b+1} - x_2^b x_4\]
and
\[x_1^{a+b+1} - x_2 x_4^b, x_1^{a+1} x_2^b - x_3 x_4^b, x_4^{b+1} - x_1^{a+1} x_3^b\]
(recall that the first three binomials generates $J$). 
In \cite{OjKa}, a complete classification of the monomial curves in $\mathbb{A}^4(\Bbbk)$ having a unique minimal system of generators is given. By \cite[Theorem 3.11]{OjKa}, one has that $I$ has a unique minimal system of generators if and only if $x_1^{a+1} x_3^b$ is not divisible by $x_1^b x_3$; equivalently $a < b-1$ as we already knew by Corollary \ref{Cor GB2}. Observe that the result on the uniqueness of the system of generators of $I$ can be deduced from \cite{KT10}, too.
\end{Example}

We end this paper by observing that, since both $J$ and $I$ are determinantal ideals by Corollaries \ref{Cor GB1} and \ref{Cor GB2}, respectively, one can conveniently adapt \cite[Section 2.1]{Philippe} to compute the minimal free resolution of $I$ and $J$ using the Eagon-Northcott complex. In particular, the $\Bbbk-$algebras $\Bbbk[x_1, \ldots, x_n]/J$ and $\Bbbk[x_1, \ldots, x_n]/I$ are Cohen-Macaulay of type $n-2$ and $n-1$, respectively (see \cite[Section 2.1]{Philippe} for further details).

\end{document}